# Denoising and Reconstruction of Nonlinear Dynamics using Truncated Reservoir Computing

### v1

@@ II.4 Tuning the Ridge Coefficient
The ridge regression coefficient ($\lambda$) becomes important when dealing with noisy data. This coefficient is inversely proportional to the level of noise contained in the data, and in the Machine Learning literature, it is often calibrated through cross-validation techniques Hastie, Tibshirani, and Friedman (2017). It should be noted that this coefficient can also be treated as a hyperparameter and calibrated via Eq. 5. However, due to the unknown nature of noise level and the wide range of potential values for this parameter, the inclusion of this parameter in the optimizing algorithm might not be feasible. Therefore, an empirical optimization algorithm is preferred, wherein the optimal value of this parameter is selected from a relatively coarse one-dimensional grid. For this purpose, a set of exponents of 10 are introduced to the cross-validation algorithm, i.e., $\{10^{-15},10^{-14},...,10^{20}\}$, and then, the best choice is specified through computing the cross-validation score. Although this approach might not be precise in terms of simultaneously optimizing all hyperparameters, it offers significant computational merits as it retains the closed-form solution of Eq. 3.
@@ II.5 Pruning and Evolution Mechanisms
The above optimization approach can be used to specify an optimal number of nodes for the RC while edges, representing the internal connectivity in the reservoir, are selected at random. Due to this arbitrary initialization, such an RC might not be in its minimal form, and there could be redundant nodes and edges, degrading the prediction performance. In this sense, it is reasonable to explore ways to reduce redundancies through truncating unnecessary nodes and edges Yadav et al. (2025); Yadav, Sinha, and Stender (2025). Attaining minimality in the size and structure of the reservoir is also important for hardware development when a digital circuit is going to be designed for implementation via a physical RC. Therefore, to remove ineffective nodes and edges, one can first train the reservoir and its hyperparameters, considering a specific number of nodes and random edges, and then perform an exhaustive search for discarding a predefined percentage of nodes and connections subject to satisfying a performance metric. The second stage can be implemented considering another optimization scheme, where the mean-squared error (MSE) calculated for the validation data set should be minimized while the pruning of nodes and edges takes place. However, this pruning strategy requires establishing a ranking mechanism across the nodes and edges so that the elimination will be started from the least influential nodes. Five measures are considered for weighing the nodes’ importance:
@@
  • Edges: Nodes with more connections may have a greater influence on the RC’s performance.
@@
  • Page-rank: This web-inspired scoring approach integrates the influence of inbound and outbound edges, as well as the quality of edges Hagberg, Schult, and Swart (2008).
The above measures can be normalized to show a score in each category between zero and one. Then, the summation of these factors can be used for selecting the nodes nominated for elimination. Based on a preselected truncation percentage, redundant nodes will be eliminated, and the RC hyperparameters will be re-calibrated. This process continues until a target performance is either achieved or the maximum number of trials is attained.
In contrast to the pruning approach, we might encounter cases in which the number of nodes appears insufficient to describe the dynamics properly, and the performance can still be improved by adding extra nodes and connections. Although increasing the number of nodes can be the option always available at the expense of extra computational costs, the performance increase may be limited since nodes and edges are initially selected at random in the classical RC framework. Additionally, the presented pruning approach is essentially a backward truncation approach, which might not cause a significant change in the performance indices. In such cases, it is reasonable to start with a simple RC with minimum dimensionality and consider an evolutionary approach for the growth of the RC toward a more complex network architecture Yadav et al. (2025). This evolution can be implemented by following the same greedy approach as above but with one difference, that is, the addition of nodes and edges shall undergo a similar post-processing optimization.
@@ III Discrete-Time Extended Kalman Filter
Kalman-type filters are well-known tools for the denoising and reconstruction of dynamic responses from incomplete noisy measurements Anderson and Moore (1979). They operate on state-space models of distributed systems and require a clear mathematical model of the system to discard noise. In Appendix A, a mathematical exposition of the EKF is provided for reference.
@@ IV.1 Lorenz Chaotic Dynamics
The Lorenz system is attractive to study in this work as it qualitatively exhibits different dynamics along the period-doubling route to chaos (see Appendix B for its governing equations and details). For generating training data, $\sigma=10$, $\rho=28$, and $\beta=8/3$ are considered. By employing a time interval of ${\Delta}t=0.005s$ and a recording duration of $T=50s$, discrete-time dynamical data sets were generated and polluted with specific levels of additive Gaussian White Noise (GWN). Across these data sets, the intensity of noise is considered variable as the noise standard deviation ranges from 0.1% to 100% of the root-mean-square (RMS) of the original signal. The noise-contaminated recordings of $x(t)$ and $y(t)$ are considered observables of a hypothetical measurement system and used as the input to the computing reservoir, whereas the clean version of all states is regarded as its output (FIG. 1). Given this specification, the generated data of $t\in[0,25s]$ acts as the training data, and that corresponding to $t\in[25s,50s]$ serves as the validation set. The reservoir hyperparameters are optimized through searching within specific intervals, including the leakage rate from $\alpha\in(0.01,1)$, the spectral radius from $\gamma\in(0.01,1)$, the input scaling from $(0.1,2)$, and the reservoir nodes connectivity from $(0.1,0.9)$.
FIGS. 1 (a) and (b) represent the inputs and outputs of the RC for reconstructing Lorenz chaotic dynamics. The results presented correspond to a reservoir computer whose hyperparameters and readout nodes are first trained, with its internal network being then truncated through the method explained above. The output of the reservoir is considerably smoother than the original noisy signals (FIG. 1). The comparison of the noisy and reproduced dynamics in phase-space (FIG. 1 (c)) shows the capability of RC in separating the dynamics of noise and responses.
@@ IV.1.1 Optimization of Nodes and Edges
This section examines to what extent the hyperparameter optimization and node removal approach can enhance the reservoir performance, particularly compared to a classical RC, in which the edges are random and only readout nodes are subject to calibration. The MSE is considered a measure of accuracy when the same validation set is used to compare different training schemes. However, due to the difference in the order of magnitude of the responses, a normalization over the mean-squared of the signal is included, yielding a dimensionless measure as follows:
[…]
Given the above measure, FIG. 2 shows improvement in the logarithm of NMSE for three training strategies considering different SNRs in the validation data set. The vanilla reservoir whose readout nodes are only trained provides the largest NMSE, whereas further optimization of the reservoir can improve the accuracy up to several orders of magnitude in some cases. More specifically, the hyperparameter tuning presented above can considerably reduce the prediction errors, and truncating nodes and edges can further improve the accuracy. However, this superior prediction performance comes at the expense of extra computational costs.
The performance of these methods is also compared with that of the EKF. As can be seen in FIG. 2, the presented truncated RC can outperform EKF at SNRs of 1 and 4 despite the fact that no explicit knowledge of the dynamics is introduced to the RC, which is contrary to the EKF that relies on the accuracy of the state-space model. Based on this result, it can be concluded that, specifically at low SNRs, the RC can offer competitive accuracy when compared with the EKF. However, it should be noted that this comparison is made while the EKF is in an advantageous position. In effect, the EKF uses physics-driven state-space models to discard noise, whereas the RC explicitly separates the desired dynamics from noise through the proposed learning protocol.
@@ IV.1.2 Characterization of Denoising Gain
In this section, we investigate when the reservoir is trained on a data set with a specific SNR and how its accuracy generalizes when tested on validation data sets with smaller or larger SNRs. For this purpose, RCs and their hyperparameters are trained using data sets contaminated with certain levels of noise. Then, the calibrated RCs are used for denoising and reconstruction of held-out data sets with different levels of noise. To compare the denoising capability in these cases, a gain ratio is introduced as follows:
[…]
@@ IV.1.3 Generalization of Denoising Performance
As demonstrated, the accuracy of the RC depends on the intensity of noise contained in the data. Similarly, it can be concluded that the accuracy might also be related to the Lorenz systematic parameters. To analyze such effects, the denoising gain is presented in FIG. 4 for validation data sets generated with the Prandtl parameter varying within $\sigma\in[0-20]$, specifically when the RC and its hyperparameters were trained on SNR = 4 and $\sigma=10$. As can be seen, the best gain is achieved in the vicinity of the selected Prandtl parameter ($\sigma=10$) used for training, considering the same noise level as the training data. Within a relatively wide neighborhood of $\sigma=10$, the gain is still significantly larger than one, although the denoising gain decreases as we deviate from the center.
Figure 4: Generalization of denoising gain in terms of variation in Prandtl parameter. (a) Denoising gain is presented when different levels of SNR and Prandtl parameters are examined for a given RC trained on a data set with $\sigma=10$ and SNR = 4. (b) Phase-space representation of Lorenz attractor for specific choices of $\sigma$. (c) Bifurcation diagram showing variation in the maximum $z(t)$. Results in sub-figure (a) are averaged over 10 randomly configured RCs to reduce the impact of randomly-initialized RCs.
It is also worth investigating how the denoising gain varies across the domain of the Prandtl parameter, considering two additional data sets for the training of the RC and its hyperparameters. One data set is generated using $\sigma=8.0$ with SNR = 1, and the other is generated using the same parameters as before, i.e., $\sigma=10.0$ but using SNR = 1. Note that all other factors are considered the same as the original data set. FIG. 5 presents the denoising gain for the validation data sets when the above settings are used for generating the training data. It is clear that the denoising gain considerably improves around $\sigma=8.0$ and particularly for SNR = 1, exhibiting a better generalization performance across the parameter domain as well. As can be seen, by including extra training data, it is possible to enhance the denoising gain and obtain more accurate computing reservoirs. This result promotes the potential of using a reservoir network trained on a few simple chaotic orbits when applied for reconstructing unseen nonlinear dynamics and highlights how the denoising gain improves as additional data is included.
@@ IV.2 AdEx Model
Three different noise processes are considered, varying in terms of the frequency content, comprising GWN, violet, and pink noise processes. The first 200 ms is treated as the training data, whereas the next 200 ms is considered as validation data. Reservoir nodes are not considered fixed and are released to vary between $[50-100]$ to represent cases in which the number of reservoir nodes is variable. Other hyperparameters are allowed to vary within the intervals specified earlier for the Lorenz attractor. According to Appendix C, simulated data is generated at 100 kHz sampling rate for a period of 400 ms and polluted with 10% RMS measurement noise (20 dB). The simulated current is considered a step function starting at 10 ms for a fixed interval of 390 ms while having 65 pA magnitude. Firing patterns are simulated through Dirac delta functions, implemented in the adaptation parameter equation. FIG. 6 shows the ground truth membrane potential and adaptation variable along with dynamics contaminated with GWN. The results of the pruned RC are also shown in FIG. 6, presenting accurate results across both training and testing data sets when compared with the ground truth. A substantial amount of noise is discarded, and the dynamic patterns are made more evident through the presented RC framework. However, the most significant mismatch between the denoised and the ground truth membrane potential prevails at the bursting intervals, where sharp peaks are not well captured. This issue can originate from the mathematical expression of ESN, which seemingly simulates smooth dynamics better.
@@
Figure 6: Reproducing noisy measurements of the AdEx model through the proposed RC framework (SNR=20 dB). The ground truth data appears in dashed blue lines, representing a neuronal behavior subjected to an initial firing pattern, followed by a subsequent adaptation. The noisy and denoised curves are shown with green and red colors. (a) Membrane potential is shown subject to fixed current input and firing patterns triggered through Dirac delta functions (b) Adaptation variable is specified in the same substance as the input current with initial firing patterns and periodic spiking behavior.
@@
The GWN process contains the same level of noise across the entire frequency band, but this assumption can be relaxed by considering colored noise. Thus, we investigate the denoising performance under violet and pink noise, where the high- and low-frequency components are dominant, respectively. The power spectral density (PSD) of violet noise varies linearly with the frequency, i.e., $S(f)\propto f$, whereas the PSD of pink noise is inversely proportional to the frequency, i.e., $S(f)\propto 1/f$. FIG. 7 compares the denoised dynamics with the noisy ones in the phase-space for the three noise processes. The denoising residual errors are also indicated in the phase-space and the frequency domain for comparison purposes. These errors are denoted by $\epsilon_{V}(t)=V(t)-\hat{V}(t)$ and $\epsilon_{w}(t)=w(t)-\hat{w}(t)$, where $\hat{V}(t)$ and $\hat{w}(t)$ represent the ground truth membrane potential and adaptation parameter, respectively.
Specifically for GWN, the noise has been reduced considerably, as shown in FIG. 7(a). The subtraction of the denoised signal and the ground truth one is shown in FIG. 7(b), highlighting the capability of RC to discard GWN up to a large extent. The PSD of the membrane potential error, $S_{\epsilon_{V}}(f)$, is shown in FIG. 7(c). Based on this figure, the reduction in the noise magnitude varies between 5 dB and 25 dB, depending on the frequency range. In effect, the residual noise is stronger in the low-frequency range of [0-5 kHz], but it reaches a weaker and stable regime around -50 dB/Hz at frequencies above 5 kHz, as shown in FIG. 7(c).
@@
Similar comparisons are made for violet noise in FIG. 7(d-f). The denoising performance is good. Notably, the residual noise appears similar to a GWN process with an almost fixed PSD equal to -55 dB/Hz. The original additive noise has a smaller magnitude at frequencies below 10 kHz, meaning that the RC produces extra noise in this particular frequency range. However, above 10 kHz, the denoising performance becomes significant, improving up to 30 dB/Hz.
@@
The pink noise primarily appears as some shift in the phase-space, as shown in FIG. 7(g). The RC mitigates noise at frequencies below 350 Hz, but above this range, it produces extra noise. Overall, the results display limited merits for RC in discarding pink noise.
@@
Figure 7: Denoising performance comparison between different additive noise processes (a-c) GWN $S(f)=S_{0}$ (d-f) Violet noise $S(f)\propto f$ (g-i) Pink noise $S(f)\propto 1/f$. The left side subplots show the denoised membrane potential and adaptation variable in the phase-space compared to the noisy ones. The middle column subplots compare residual noise with the originally added noise in the phase-space. The right side subplots show the averaged PSD of the residual and original noise processes.

### v2

@@ II.4 Tuning the Ridge Coefficient
The ridge regression coefficient ($\lambda$) becomes important when dealing with noisy data. This coefficient is inversely proportional to the level of noise contained in the data, and in the machine learning literature, it is often calibrated through cross-validation techniques Hastie, Tibshirani, and Friedman (2017). It should be noted that this coefficient can also be treated as a hyperparameter and calibrated via Eq. 5. However, due to the unknown nature of noise level and the wide range of potential values for this parameter, the inclusion of this parameter in the optimization algorithm might not be feasible. Therefore, an empirical optimization algorithm is preferred, wherein the optimal value of this parameter is selected from a relatively coarse one-dimensional grid. For this purpose, a set of exponents of ten is introduced to the cross-validation algorithm, i.e., $\{10^{-15},10^{-14},...,10^{20}\}$, and then, the best choice is specified through computing the cross-validation score. Although this approach might not be precise in terms of simultaneously optimizing all hyperparameters, it offers significant computational merits as it retains the closed-form solution of Eq. 3.
@@ II.5 Pruning and Evolution Mechanisms
The above optimization approach can be used to specify an optimal number of nodes for the RC while edges, representing the internal connectivity in the reservoir, are selected at random. Due to this arbitrary initialization, such an RC might not be in its minimal form, and there could be redundant nodes and edges, degrading the prediction performance. In this sense, it is reasonable to explore ways to reduce redundancies through truncating unnecessary nodes and edges Yadav et al. (2025); Yadav, Sinha, and Stender (2025). Attaining minimality in the size and structure of the reservoir is also important for hardware development when a digital circuit is going to be designed for implementation via a physical RC Kent, Barbosa, and Gauthier (2024). Therefore, to remove ineffective nodes and edges, one can first train the reservoir and its hyperparameters, considering a specific number of nodes and random edges, and then perform an exhaustive search for discarding a predefined percentage of nodes and connections, subject to satisfying a performance metric. The second stage can be implemented considering another optimization scheme, where the mean-squared error (MSE) calculated for the validation data set should be minimized while the pruning of nodes and edges takes place. However, this pruning strategy requires establishing a ranking mechanism across the nodes and edges so that the elimination can start from the least influential nodes. Five measures are considered for weighing the nodes’ importance:
@@
  • Number of edges: Nodes with more connections may have a greater influence on the RC’s performance.
@@
  • Page-rank: This web-inspired scoring approach integrates the influence of inbound and outbound edges, as well as the quality of edges as measured in terms of their relative weight, frequency of interaction, and contextual relevance within the network.Hagberg, Schult, and Swart (2008).
The above measures can be normalized to show a score in each category between zero and one. Then, the summation of these factors can be used for selecting the nodes nominated for elimination. Based on a preselected truncation percentage, redundant nodes will be eliminated, and the RC hyperparameters will be recalibrated. This process continues until a target performance is either achieved or the maximum number of trials is attained. Here, we hypothesize the usefulness of such metrics based on our intuition, but choosing other metrics for pruning nodes and edges is also possible. However, the decision whether to remove or retain the nominated nodes or edges must be made based on the performance.
In contrast to the pruning approach, we might encounter cases in which the number of nodes appears insufficient to describe the dynamics properly, and the performance can still be improved by adding extra nodes and connections. Although increasing the number of nodes can be an option always available at the expense of extra computational costs, the performance increase may be limited since nodes and edges are initially selected at random in the classical RC framework. Additionally, the presented pruning approach is essentially a backward truncation approach, which might not cause a significant change in the performance indices. In such cases, it is reasonable to start with a simple RC with minimum dimensionality and consider an evolutionary approach for the growth of the RC toward a more complex network architecture Yadav et al. (2025). This evolution can be implemented by following the same greedy approach as above, but with one difference, that is, the addition of nodes and edges shall undergo a similar post-processing optimization.
@@ III Discrete-Time Extended Kalman Filter
To provide a competitive baseline, we consider a qualitatively different filtering technique based on explicit modeling of the system dynamics. A method widely applied to nonlinear systems is EKF, a well-known tool for the denoising and reconstruction of dynamic responses from incomplete noisy measurements Anderson and Moore (1979). EKF operates on state-space models of distributed systems and requires a clear mathematical model of the system to discard noise. In Appendix A, a mathematical exposition of the EKF is provided for reference.
@@ IV.1 Lorenz Chaotic Dynamics
The Lorenz system is attractive to study in this work as it qualitatively exhibits different dynamics along the period-doubling route to chaos (see Appendix B for its governing equations and details). For generating training data, $\sigma=10$, $\rho=28$, and $\beta=8/3$ are considered. By employing a time interval of ${\Delta}t=0.005s$ and a recording duration of $T=50s$, discrete-time dynamical data sets were generated considering initial conditions $x_{0}=y_{0}=z_{0}=1$ and polluted with specific levels of additive Gaussian White Noise (GWN). Across these data sets, the intensity of noise is considered variable as the noise standard deviation ranges from 0.25% (52 dB) to 100% (0 dB) of the root-mean-square (RMS) of the original signal’s components. The noise-contaminated recordings of $x(t)$ and $y(t)$ are considered observables of a hypothetical measurement system and used as the input to the reservoir computer, whereas the clean version of all states is regarded as its output during training (FIG. 1). Hence, the RC must learn both denoising the inputs and reconstructing a third state. Given this specification, the generated data of $t\in[0,25s]$ acts as the training data, and that corresponding to $t\in[25s,50s]$ serves as the validation set. The reservoir hyperparameters are optimized through searching within specific intervals, including the leakage rate from $\alpha\in(0.01,1)$, the spectral radius from $\gamma\in(0.01,1)$, the input scaling from $\zeta\in(0.1,2)$, and the internal connectivity of reservoir nodes from $p\in(0.1,0.9)$.
FIGS. 1 (a) and (b) represent the inputs and outputs of the RC for reconstructing Lorenz chaotic dynamics. The results presented correspond to a reservoir computer whose hyperparameters and readout nodes are first trained, with its internal network being then truncated through the method explained above. The output of the reservoir is considerably smoother than the original noisy signals (FIG. 1). The comparison of the noisy and reproduced dynamics in phase space (FIG. 1 (c)) shows the capability of RC in separating the dynamics of noise and responses.
@@ IV.1.1 Optimization of Nodes and Edges
This section examines how the hyperparameter optimization and node removal approach can enhance the reservoir performance, particularly compared to a classical RC, in which the edges are random and only readout nodes are subject to calibration. The MSE is considered a measure of accuracy when the same validation set is used to compare different training schemes. However, due to the difference in the order of magnitude of the responses, a normalization over the mean-squared of the signal is included, yielding a dimensionless measure as follows:
$\mathrm{NMSE}=\frac{\|\mathbf{\hat{Y}_{\it{n}}-\mathbf{R_{\it{n}}}\mathbf{\hat{W}}_{\mathrm{out}}}\|_{2}^{2}}{\|\mathbf{\hat{Y}_{\it{n}}}\|_{2}^{2}}$ (6)
Given the above measure, FIG. 2 shows improvement in the logarithm of NMSE for three training strategies considering different SNRs in the validation data set. Note that SNR is defined here as the RMS of the signal over the RMS of noise. The vanilla reservoir, whose readout nodes are only trained, provides the largest NMSE, whereas further optimization of the reservoir can improve the accuracy up to several orders of magnitude in some cases. More specifically, the hyperparameter tuning presented above can considerably reduce the prediction errors, and truncating nodes and edges can further improve the accuracy. However, this superior prediction performance comes at the expense of extra computational costs.
The performance of these methods is also compared with that of the EKF. As can be seen in FIG. 2, the presented truncated RC can outperform EKF at SNRs of 1 and 4 although no explicit knowledge of the dynamics is introduced to the RC, which is contrary to the EKF that relies on the accuracy of the state-space model. Based on this result, it can be concluded that, specifically at low SNRs, the RC can offer competitive accuracy when compared with the EKF. However, it should be noted that this comparison is made while the EKF is in an advantageous position. In effect, the EKF uses physics-driven state-space models to discard noise, whereas the RC explicitly separates the desired dynamics from noise through the proposed learning protocol.
@@ IV.1.2 Characterization of Denoising Gain
In this section, we investigate when the reservoir is trained on a dataset with a specific SNR and how its accuracy generalizes when tested on validation datasets with smaller or larger SNRs. For this purpose, RCs and their hyperparameters are trained using datasets contaminated with varying levels of noise. Then, the calibrated RCs are used for denoising and reconstruction of held-out data sets with different levels of noise. To compare the denoising capability in these cases, a gain ratio is introduced as follows:
$\mathrm{Denoising\ Gain}=\frac{\mathrm{SNR}_{\mathrm{Reconstructed}}}{\mathrm{SNR}_{\mathrm{Test}}}$ (7)
where ${\mathrm{SNR}_{\mathrm{Reconstructed}}}$ represents the amount of noise present in the predictions made by the RC, and ${\mathrm{SNR}_{\mathrm{Test}}}$ represents the amount of noise in the test data set. When this index is greater than 1.0, it can be concluded that the RC can discard the noise, whereas values equal to or smaller than 1.0 represent cases where the RC adversely increases the amount of noise in the reconstructed data. The color-map matrix in FIG. 3 visualizes the denoising gain, where the green and red colors show positive and negative denoising performance, respectively. Values on the minor diagonal of the matrix show cases where the training and testing data sets have had the same level of noise intensity. The best denoising gain is achieved when the training and testing data sets correspond to the same level of noise. Additionally, the denoising gain is better for large SNRs. For $\mathrm{SNR}_{\mathrm{Train}}=100\ \textnormal{and}\ 400$ that indicates the level noise in the training set, the RC is rather unsuccessful in denoising the signals, although it still has merits in terms of reconstructing the unobserved dynamics. Additionally, the color map is completely non-symmetrical, which implies that, for a specific level of noise contained in the test dataset, those RCs trained on smaller noise perform relatively better.
@@ IV.1.3 Generalization of Denoising Performance
As demonstrated, the accuracy of the RC depends on the intensity of noise contained in the data. Similarly, it can be concluded that the accuracy might also be related to the Lorenz qualitative dynamics. To analyze such effects, the denoising gain is presented in FIG. 4 for validation data sets generated with the Prandtl parameter varying within $\sigma\in[0-20]$, specifically when the RC and its hyperparameters were trained on SNR = 4 and $\sigma=10$. As can be seen, the best gain is achieved in the vicinity of the selected Prandtl parameter ($\sigma=10$) used for training, considering the same noise level as the training data. Within a relatively wide neighborhood of $\sigma=10$, the gain is still significantly larger than one, although the denoising gain decreases as we deviate from the center. However, there is a pattern change for values of $\sigma<4.0$, where the denoising gain increases. This observation is attributed to the gradual change in the nonlinear dynamics in the vicinity of $\sigma=4.0$, as demonstrated in the bifurcation diagram of FIG. 4(c). This figure highlights that the denoising gain also depends on the nonlinear dynamics considered for training the RC, and it will not necessarily show monotonous behavior as we deviate from the neighborhood of the parameter used for training the RC.
Figure 4: Generalization of denoising gain in terms of variation in Prandtl parameter. (a) Denoising gain is presented when different levels of SNR and Prandtl parameters are examined for a given RC trained on a data set with $\sigma=10$ and SNR = 4. (b) phase space representation of Lorenz attractor for specific choices of $\sigma$. (c) Bifurcation diagram showing variation in the maximum $x(t)$. Results in sub-figure (a) are averaged over 10 randomly configured RCs to reduce the impact of randomly-initialized RCs.
It is also worth investigating how the denoising gain varies across the domain of the Prandtl parameter, considering two additional data sets for the training of the RC and its hyperparameters. One data set is generated using $\sigma=8$ with SNR = 1, and the other is generated using the same parameters as before, i.e., $\sigma=10$ but using SNR = 1. Note that all other factors are considered the same as the original data set. FIG. 5 presents the denoising gain for the validation data sets when the above settings are used for generating the training data. It is clear that the denoising gain considerably improves around $\sigma=8$ and particularly for SNR = 1, exhibiting a better generalization performance across the parameter domain as well. As can be seen, by including extra training data, it is possible to enhance the denoising gain and obtain more accurate computing reservoirs. This result promotes the potential of using a reservoir network trained on a few simple chaotic orbits when applied for reconstructing unseen nonlinear dynamics and highlights how the denoising gain improves as additional data is included.
@@ IV.2 AdEx Model
Three different noise processes are considered, varying in terms of the frequency content, comprising GWN, violet, and pink noise processes. The first 200 ms is treated as the training data, whereas the next 200 ms is considered as validation data. Reservoir nodes are not considered fixed and are released to vary between $[50-100]$ to represent cases in which the number of reservoir nodes is variable. Other hyperparameters are allowed to vary within the intervals specified earlier for the Lorenz attractor. According to Appendix C, simulated data is generated at 100 kHz sampling rate for a period of 400 ms and polluted with 10% RMS measurement noise (20 dB). The simulated current is considered a step function starting at 10 ms for a fixed interval of 390 ms, with a magnitude of 65 pA. Firing patterns are simulated through Dirac delta functions, implemented in the adaptation parameter equation. FIG. 6 shows the ground truth membrane potential and adaptation variable along with dynamics contaminated with GWN. The results of the pruned RC are also shown in FIG. 6, presenting accurate results across both training and testing data sets when compared with the ground truth. A substantial amount of noise is discarded, and the dynamic patterns are made more evident through the presented RC framework. However, the most significant mismatch between the denoised and the ground truth membrane potential prevails at the bursting intervals, where sharp peaks are not well captured. This issue can originate from the mathematical expression of ESN, which seemingly simulates smooth dynamics better.
@@
Figure 6: Reproducing noisy measurements of the AdEx model through the proposed RC framework (SNR=20 dB). The ground truth data appears in dashed blue lines, representing a neuronal behavior subjected to an initial firing pattern, followed by a subsequent adaptation. The noisy and denoised curves are shown in green and red colors. (a) Membrane potential is shown subject to fixed current input and firing patterns triggered through Dirac delta functions (b) Adaptation variable is specified in the same substance as the input current with initial firing patterns and periodic spiking behavior.
@@
The GWN process contains the same level of noise across the entire frequency band, but this assumption can be relaxed by considering colored noise. Thus, we investigate the denoising performance under violet and pink noise, where the high- and low-frequency components are dominant, respectively. The power spectral density (PSD) of violet noise varies linearly with the frequency, i.e., $S(f)\propto f$, whereas the PSD of pink noise is inversely proportional to the frequency, i.e., $S(f)\propto 1/f$. FIG. 7 compares the denoised dynamics with the noisy ones in the phase space for the three noise processes. The denoising residual errors are also indicated in the phase space and the frequency domain for comparison purposes. These errors are denoted by $\epsilon_{V}(t)=V(t)-\hat{V}(t)$ and $\epsilon_{w}(t)=w(t)-\hat{w}(t)$, where $\hat{V}(t)$ and $\hat{w}(t)$ represent the ground truth membrane potential and adaptation parameter, respectively.
Specifically for GWN, the noise has been reduced considerably, as shown in FIG. 7(a). The subtraction of the denoised signal and the ground truth one is shown in FIG. 7(b), highlighting the capability of RC to discard GWN to a large extent. As can be seen, the errors are more pronounced along the horizontal axis. This observation can be justified based on the time-domain comparison shown in FIG. 6, where the voltage errors $\epsilon_{V}(t)$ are large around the sharp peaks, whereas the adaptation parameter errors $\epsilon_{w}(t)$ are small. Appendix D provides further comparison between the PSD curves of the noisy and denoised signals in comparison with the ground truth PSD.
@@
The PSD of the membrane potential error, $S_{\epsilon_{V}}(f)$, is shown in FIG. 7(c). Based on this figure, the reduction in the noise magnitude varies between 5 dB and 25 dB, depending on the frequency range. In effect, the residual noise is stronger in the low-frequency range of [0-5 kHz], but it reaches a weaker and stable regime around -50 dB/Hz at frequencies above 5 kHz, as shown in FIG. 7(c). Similar comparisons are made for violet noise in FIG. 7(d-f). The denoising performance is good. Notably, the residual noise appears similar to a GWN process with an almost fixed PSD equal to -55 dB/Hz. The original additive noise has a smaller magnitude at frequencies below 10 kHz, meaning that the RC produces extra noise in this particular frequency range. However, above 10 kHz, the denoising performance becomes significant, improving up to 30 dB/Hz.
@@
The pink noise primarily appears as some shift in the phase space, as shown in FIG. 7(g). The RC mitigates noise at frequencies below 350 Hz, but above this range, it produces extra noise. Overall, the results display limited merits for RC in discarding pink noise.
@@
Figure 7: Denoising performance comparison between different additive noise processes (a-c) GWN $S(f)=S_{0}$ (d-f) Violet noise $S(f)\propto f$ (g-i) Pink noise $S(f)\propto 1/f$. The left side subplots show the denoised membrane potential and adaptation variable in the phase space compared to the noisy ones. The middle column subplots compare residual noise with the originally added noise in the phase space. The right side subplots show the averaged PSD of the residual and original noise processes.